OpenStudy (anonymous):

Solve the following system of equations. 3x + 3y + 2z = -5 x + y + 3z = -4 2x + 2y + 5z = -7

OpenStudy (anonymous):

do your own work feather

OpenStudy (anonymous):

Helpppppppppppppp

OpenStudy (unklerhaukus):

3x + 3y + 2z = -5 i x + y + 3z = -4 ii 2x + 2y + 5z = -7 iii ii+iii 3x + 3y + 8z = -11

OpenStudy (anonymous):

How what? im really bad at math

OpenStudy (unklerhaukus):

ill show you what i did ive added the left hand sides together and also the right hands sides together ii+iii (x + y + 3z)+[2x + 2y + 5z]= (-4) +[-7] x +2x + y+2y+ 3z+5z= -4 +-7 3x + 3y + 8z = -11

OpenStudy (anonymous):

A. (-4, -3, 1) B. (-4, 3, -1) C. (4, -3, -1) D. (-4, -3, -1)

OpenStudy (anonymous):

i have too get an answer like that

OpenStudy (unklerhaukus):

does what i have done so far make sense to you?

OpenStudy (anonymous):

yes ! (:

OpenStudy (unklerhaukus):

lets call our result iv, {just to give it a name} 3x + 3y + 8z = -11 iv

OpenStudy (unklerhaukus):

now what do you get when you takeaway i from iv iv - i (3x + 3y + 8z)-[3x + 3y + 2z] = (-11)-[-5] can you simplify this @alexislovier?

OpenStudy (anonymous):

i have no idea !

OpenStudy (unklerhaukus):

its a very similar to the first step

OpenStudy (anonymous):

okayy

OpenStudy (unklerhaukus):

try it